# DiffuseDef: Improved Robustness to Adversarial Attacks

### v1

@@ Abstract
Pretrained language models have significantly advanced performance across various natural language processing tasks. However, adversarial attacks continue to pose a critical challenge to system built using these models, as they can be exploited with carefully crafted adversarial texts. Inspired by the ability of diffusion models to predict and reduce noise in computer vision, we propose a novel and flexible adversarial defense method for language classification tasks, DiffuseDef111https://github.com/Nickeilf/DiffuseDef, which incorporates a diffusion layer as a denoiser between the encoder and the classifier. During inference, the adversarial hidden state is first combined with sampled noise, then denoised iteratively and finally ensembled to produce a robust text representation. By integrating adversarial training, denoising, and ensembling techniques, we show that DiffuseDef improves over different existing adversarial defense methods and achieves state-of-the-art performance against common adversarial attacks.
DiffuseDef: Improved Robustness to Adversarial Attacks
Zhenhao Li, Marek Rei, Lucia Specia Language and Multimodal AI (LAMA) Lab, Imperial College London {zhenhao.li18, marek.rei, l.specia}@imperial.ac.uk
@@ 1 Introduction
Pretrained language models (PLM) have significantly advanced the performance of various natural language processing (NLP) tasks. Despite such improvements, current NLP systems remain susceptible to adversarial attacks where carefully crafted text perturbations can lead to incorrect model outputs (Alzantot et al., 2018; Jin et al., 2020; Li et al., 2020). In order to improve robustness to adversarial attacks, various defense methods have been proposed, such as adversarial training (Zhu et al., 2020; Si et al., 2021; Zhou et al., 2021; Xi et al., 2022), text denoising (Nguyen Minh and Luu, 2022; Wang et al., 2023), ensembling (Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), etc. However, existing defense methods either assume the test-time perturbation/attack set is similar to that used in training (Li et al., 2021), or are limited to specific architectures (Xi et al., 2022), or at inference time require large computational cost, thereby limiting their practical applicability.
@@
Diffusion models are commonly used in computer vision (CV) to generate high-quality images by predicting and removing noise from a sampled noisy image. Therefore, they can be adopted to remove noise from adversarial images and thus improve robustness to attacks (Nie et al., 2022). However, in NLP very limited research has investigated adversarial defense with diffusion models due to the discrete and contextual nature of text data. Li et al. (2023) adopt the idea of iterative denoising and reconstruct adversarial texts from masked texts, while Yuan et al. (2024) use a diffusion model as a classifier and perform reverse diffusion steps on the label vector, conditioning on the input text. Inspired by the general noise prediction and reduction capability of diffusion models, we propose DiffuseDef, a novel adversarial defense method which employs diffusion training to denoise hidden representations of adversarial texts. Unlike Li et al. (2023) and Yuan et al. (2024) which apply diffusion on texts or labels, DiffuseDef directly removes noise to the hidden states, providing a more effective and robust text representation to defend against adversarial texts. Compared to diffusion-based defense in CV (Nie et al., 2022), DiffuseDef further enhances robustness with ensembling and improves efficiency with fewer diffusion steps.
@@
DiffuseDef combines adversarial training with diffusion training, where the diffusion layer is trained to predict randomly sampled noise at a given timestep. During inference, the diffusion layer serves as a denoiser, iteratively removing noise from adversarial hidden states to yield a robust hidden representation. Moreover, we adopt the ensembling strategy by first adding random noise to text hidden states to create multiple variants then denoising them via the diffusion layer. The model output is made by averaging all denoised hidden states. Since ensembling happens solely at the diffusion layer, DiffuseDef is more efficient than traditional ensembling-based methods (Ye et al., 2020; Zeng et al., 2023), which require a full forward pass through all model parameters.
@@
Through systematic experimentation, we demonstrate that DiffuseDef outperforms strong defense methods and is able to defend against multiple types of adversarial attacks, while preserving performance on clean texts. Our analysis also reveals that the ensembling diffused representations provides a stronger defense against finding vulnerable words to attack and can reduce the distance in latent space between adversarial texts and their clean text counterpart.
Our contributions can be summarised as follows:
@@
DiffuseDef outperforms existing adversarial methods and achieves state-of-the-art performance against prevalent adversarial attacks.
@@ 2.1 Textual Adversarial Attacks
Textual adversarial attacks focus on constructing adversarial examples from an original text that maximise the likelihood of incorrect predictions by a neural network. These attacks require adversarial examples to be perceptually similar to the original text, which is typically achieved by introducing subtle perturbations to the original text, such as character swapping (Gao et al., 2018; Ebrahimi et al., 2018), synonym-substitutions (Ren et al., 2019; Yoo and Qi, 2021), and paraphrasing (Gan and Ng, 2019; Huang and Chang, 2021). Taking the text classification task as an example, given a classifier that maps an input sequence of words to its designated label , the goal of the attack model is to construct an adversarial example to fool the classifier, where is a subtle adversarial perturbation constrained by . The adversarial example is considered a successful attack if it leads to an incorrect prediction . The attacker can iteratively generate multiple adversarial examples and query the classifier to obtain a successful attack, whereas the classifier must consistently return the correct prediction within a specified number of query attempts to be considered robust.
Common textual adversarial attack methods adopt a two-stage process to construct effective adversarial examples: word importance ranking and word substitution. In the first stage, words or subwords are ranked based on their influence on the model’s prediction. This is measured by leveraging either gradient information (Liu et al., 2022) or changes in prediction probabilities when words are removed (Jin et al., 2020) or masked (Ren et al., 2019; Li et al., 2020). In the second stage, candidate words are substituted with synonyms (Zang et al., 2020), perturbed variants (Gao et al., 2018), or outputs from masked language models (Garg and Ramakrishnan, 2020; Li et al., 2020). The substitution process is guided by various constraints to ensure the adversarial example remains natural and semantically equivalent to the original text. Common constraints include thresholding the similarity between the replacement word embedding and the substituted word embedding, or ensuring the semantic similarity between sentence vectors modeled from Universal Sentence Encoder (Cer et al., 2018). Despite these constraints, current textual adversarial attacks still pose significant challenges to NLP models (Liu et al., 2022; Xu et al., 2021; Yuan et al., 2023), highlighting the necessity for defense methods for better adversarial robustness.
@@ 2.2 Adversarial Defense Methods
To mitigate the performance degradation caused by adversarial attacks, various adversarial defense methods have been developed. They can be grouped into three categories: training-based, ensembling-based, and denoising-based methods. Adversarial training improves the robustness of the model to adversarial examples through strategies like data augmentation (Si et al., 2021) and adversarial regularisation (Madry et al., 2018; Zhu et al., 2020; Wang et al., 2021; Xi et al., 2022; Gao et al., 2023). However, adversarial training methods are limited as they assumes similar train-test adversarial examples, and thus tend to overfit to specific types of adversarial attacks. Ensembling-based methods generate multiple variants of the input text at inference time and ensemble model predictions over all the variants (Ye et al., 2020; Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), but they can be inefficient given that model predictions are needed on every ensemble, increasing the inference time with the number of ensembles. More recently, denoising-based methods have been proposed to improve adversarial robustness by mapping the vector representation of the adversarial text to another point in the latent space that is close to the clean text (Nguyen Minh and Luu, 2022; Wang et al., 2023; Moon et al., 2023; Yuan et al., 2024). The denoised representation makes it more difficult to find vulnerable words to attack, thus improving adversarial robustness (Wang et al., 2023). Nevertheless, denoising might lead to very different representations of clean text and adversarial text, therefore changing the semantic meanings.
@@ 3.1 Training
The proposed diffusion defense model consists of a pretrained encoder for feature extraction, a transformer-based diffusion layer for noise prediction and reduction, and a classifier layer for output generation. The training process is split into two stages: adversarial training and diffusion training (Figure 1). The adversarial training stage employs any neural network-based adversarial training methods like FreeLB++ (Li et al., 2021) and RSMI (Moon et al., 2023), which optimise the encoder and classifier for robustness by perturbing the latent representation of the text input.
In the diffusion training stage, only the diffusion layer is trained to predict random noise added to the clean text hidden state at different timesteps, enabling it to denoise the adversarial hidden state at inference time. The pretrained encoder, however is frozen during this stage. Since the pretrained encoder is only used for feature extraction, the diffusion training method is compatible with any neural network-based adversarial training method.
Given an input sequence of tokens , the pretrained encoder extracts the hidden state . A random Gaussian noise is sampled to perturb hidden state . Sohl-Dickstein et al. (2015) define the forward diffusion process as a Markov Chain where at each timestep a Gaussian noise is sampled and added to the previous latent feature: , where , is the noisy hidden state at step and is a pre-calculated variance schedule changing with . As shown by Ho et al. (2020), this equation can be reformulated to calculate directly from by defining and , thus
At each training step, a random forward diffusion timestep is sampled from a uniform distribution. Therefore, the noisy hidden state is created from , , and . The diffusion layer consists of a time embedding and a transformer layer. The time embedding receives the diffusion timestep as input and produces an embedding , which is added to as input for the transformer layer. Finally, the transformer layer outputs the predicted noise , and mean square error is used to compute the loss between predicted noise and actual sampled noise .
@@ 3.2 Inference
Leveraging the diffusion layer’s ability to predict noise at a given timestep , we utilise it as a denoiser during inference by iteratively performing the reverse diffusion steps, which sample from to produce the denoised hidden state
@@
Inference in DiffuseDef combines a one-step noising, a multi-step denoising, and an ensembling step. After the pretrained encoder extracts its hidden state , a set of Gaussian noise vectors are sampled to perform a single forward diffusion step. These noise vectors are then added to the hidden state following equation 1, resulting in a set of noisy hidden states , where denotes the number of denoising steps. The noisy hidden states are subsequently denoised through reverse diffusion steps, where noise is predicted by the diffusion layer and subtracted from the previous noisy hidden states. Unlike Ho et al. (2020) where the reverse diffusion step starts with pure noise sampled from standard normal distribution, we assume the noisy hidden state is already an intermediate state in the reverse diffusion steps. This allows us to use a smaller number of than the training timestep to prevent the denoised hidden states from diverging substantially from the initial hidden state . This sequence of denoising steps creates the final denoised hidden states , which are averaged and used by the classifier to output the final predicted label. This process is summarised in Algorithm 1.
@@ 4 Experiments
Datasets
We focus on two common NLP tasks in our experiments: topic classification and natural language inference (NLI). In the text classification task, we compare our method with other defense algorithms on two standard datasets for adversarial defense: AG News (Zhang et al., 2015a) and IMDB (Maas et al., 2011a) datasets. In the NLI task, we perform an ablation analysis with the Question-answering NLI (QNLI) dataset (Wang et al., 2018). We randomly split AGNews, IMDB, and QNLI datasets into train, validation, and test splits.
Evaluation
Following previous work on adversarial defense, we use three benchmarking attack methods to evaluate the robustness of DiffuseDef: TextFooler (TF) (Jin et al., 2020), TextBugger (TB) (Li et al., 2019), and Bert-Attack (BA) (Li et al., 2020). The three attack methods create adversarial attacks in different granularities: character-level perturbation (TextBugger), word substitution (TextFooler), and subword substitution (BertAttack). Regarding evaluation metrics, we measure the clean accuracy (Clean%) on the test set, the accuracy under attack (AUA%), and the number of adversarial queries (#Query) needed for a successful attack. Higher scores on the three metrics denote a better robustness performance of a defense method. The accuracy on clean data is measured across the entire test set. The accuracy under attack and number of queries, due to the lengthy attacking process, is measured on a randomly sampled subset of 1000 examples from the test set. We use the TextAttack library as the adversarial evaluation framework. To ensure a fair comparison and high-quality adversarial examples, we follow the same evaluation constraints as in Li et al. (2021). The evaluation metrics are averaged based on experiments run with 5 random seeds.
@@ 4.1 Comparison to SOTA
We compare our proposed method with state-of-the-art adversarial defense approaches, trained using both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as backbones: Fine-tune: Fine-tuning pretrained models on downstream task with no defense method applied222”Fine-tune” is a baseline approach used to illustrate the effect of adversarial attacks.. InfoBERT (Wang et al., 2021): Applying mutual-information-based regularizers during fine-tuning of pretrained models to improve robustness. FreeLB++ (Li et al., 2021): An adversarial training method improving on FreeLB(Zhu et al., 2020), which adds adversarial perturbations to word embedding during fine-tuning. EarlyRobust333We only run EarlyRobust with BERT as its implementation with RoBERTa has not been released. (Xi et al., 2022): Extracting early-bird subnetworks and pruning pretrained models for efficient adversarial training. RSMI (Moon et al., 2023): A two-stage training method that combines randomised smoothing and masked inference to improve adversarial robustness.

### v2

@@ Abstract
Pretrained language models have significantly advanced performance across various natural language processing tasks. However, adversarial attacks continue to pose a critical challenge to systems built using these models, as they can be exploited with carefully crafted adversarial texts. Inspired by the ability of diffusion models to predict and reduce noise in computer vision, we propose a novel and flexible adversarial defense method for language classification tasks, DiffuseDef 111Codes available at: https://github.com/Nickeilf/DiffuseDef, which incorporates a diffusion layer as a denoiser between the encoder and the classifier. The diffusion layer is trained on top of the existing classifier, ensuring seamless integration with any model in a plug-and-play manner. During inference, the adversarial hidden state is first combined with sampled noise, then denoised iteratively and finally ensembled to produce a robust text representation. By integrating adversarial training, denoising, and ensembling techniques, we show that DiffuseDef improves over existing adversarial defense methods and achieves state-of-the-art performance against common black-box and white-box adversarial attacks.
DiffuseDef: Improved Robustness to Adversarial Attacks via Iterative Denoising
Zhenhao Li, Huichi Zhou, Marek Rei, Lucia Specia Language and Multimodal AI (LAMA) Lab, Imperial College London {zhenhao.li18, h.zhou24, marek.rei, l.specia}@imperial.ac.uk
@@ 1 Introduction
Pretrained language models (PLM) have significantly advanced the performance of various natural language processing (NLP) tasks. Despite such improvements, current NLP systems remain susceptible to adversarial attacks where carefully crafted text perturbations can lead to incorrect model outputs (Alzantot et al., 2018; Jin et al., 2020; Li et al., 2020). In order to improve robustness to adversarial attacks, various defense methods have been proposed, such as adversarial training (Zhu et al., 2020; Si et al., 2021; Zhou et al., 2021; Xi et al., 2022), text denoising (Nguyen Minh and Luu, 2022; Wang et al., 2023), ensembling (Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), etc. However, existing defense methods either assume the test-time perturbation/attack set is similar to that used in training (Li et al., 2021), or are limited to specific architectures (Xi et al., 2022), or at inference time require large computational costs, thereby limiting their practical applicability.
@@
Diffusion models are commonly used in computer vision (CV) to generate high-quality images by predicting and removing noise from a sampled noisy image. Therefore, they can be adopted to remove noise from adversarial images and thus improve robustness to attacks (Im et al., 2016; Nie et al., 2022; Gulsen et al., 2024). However, in NLP very limited research has investigated adversarial defense with diffusion models due to the discrete and contextual nature of text data. Li et al. (2023) adopt the idea of iterative denoising and reconstruct adversarial texts from masked texts, while Yuan et al. (2024) use a diffusion model as a classifier and perform reverse diffusion steps on the label vector, conditioning on the input text. Inspired by the general noise prediction and reduction capability of diffusion models, we propose DiffuseDef, a novel adversarial defense method which employs diffusion training to denoise hidden representations of adversarial texts. Unlike Li et al. (2023) and Yuan et al. (2024), which apply diffusion on texts or labels, DiffuseDef directly removes noise from the hidden states, providing a more effective and robust text representation to defend against adversarial texts. Different from diffusion-based defenses in CV (Nie et al., 2022; Gulsen et al., 2024) that apply purification to input images, DiffuseDef only performs denoising at the final hidden layer, resulting in an improvement in efficiency. This improvement is further enhanced when combined with ensembling, as it eliminates the need for a forward pass through all classifier parameters for each ensemble.
@@
DiffuseDef combines adversarial training with diffusion training, where the diffusion layer is trained to predict randomly sampled noise at a given timestep. During inference, the diffusion layer serves as a denoiser, iteratively removing noise from adversarial hidden states to yield a robust hidden representation. Moreover, we adopt the ensembling strategy by first adding random noise to the text hidden states to create multiple variants, then denoising them via the diffusion layer. The model output is made by averaging all denoised hidden states. Since ensembling happens solely at the diffusion layer, DiffuseDef is more efficient than traditional ensembling-based methods (Ye et al., 2020; Zeng et al., 2023), which require a full forward pass through all model parameters.
@@
Through systematic experimentation, we demonstrate that DiffuseDef outperforms strong defense methods and is able to defend against multiple types of black-box and white-box adversarial attacks, while preserving performance on clean texts. Our analysis also reveals that the ensembling diffused representation provides a stronger defense against finding vulnerable words to attack and can reduce the distance in latent space between adversarial texts and their clean text counterparts.
Our contributions can be summarized as follows:
@@
DiffuseDef outperforms existing adversarial methods and achieves state-of-the-art performance against prevalent black-box and white-box adversarial attacks.
@@ 2.1 Textual Adversarial Attacks
Textual adversarial attacks focus on constructing adversarial examples from an original text that maximize the likelihood of incorrect predictions by a neural network. These attacks require adversarial examples to be perceptually similar to the original text, which is typically achieved by introducing subtle perturbations to the original text, such as character swapping (Gao et al., 2018; Ebrahimi et al., 2018), synonym-substitutions (Ren et al., 2019; Yoo and Qi, 2021; Liu et al., 2023), and paraphrasing (Gan and Ng, 2019; Huang and Chang, 2021). Taking the text classification task as an example, given a classifier that maps an input sequence of words to its designated label , the goal of the attack model is to construct an adversarial example to fool the classifier, where is a subtle adversarial perturbation constrained by . The adversarial example is considered a successful attack if it leads to an incorrect prediction . The attacker can iteratively generate multiple adversarial examples and query the classifier to obtain a successful attack, whereas the classifier must consistently return the correct prediction within a specified number of query attempts to be considered robust.
Common textual adversarial attack methods adopt a two-stage process to construct effective adversarial examples: word importance ranking and word substitution. In the first stage, words or subwords are ranked based on their influence on the model’s prediction. This is measured by leveraging either gradient information (Liu et al., 2022) or changes in prediction probabilities when words are removed (Jin et al., 2020) or masked (Ren et al., 2019; Li et al., 2020). In the second stage, candidate words are substituted with synonyms (Zang et al., 2020), perturbed variants (Gao et al., 2018), or outputs from masked language models (Garg and Ramakrishnan, 2020; Li et al., 2020). The substitution process is guided by various constraints to ensure the adversarial example remains natural and semantically equivalent to the original text. Common constraints include thresholding the similarity between the replacement word embedding and the substituted word embedding, or ensuring the semantic similarity (Zhou et al., 2024) between sentence vectors modeled from Universal Sentence Encoder (Cer et al., 2018). Despite these constraints, current textual adversarial attacks still pose significant challenges to NLP models (Liu et al., 2022; Xu et al., 2021; Yuan et al., 2023), highlighting the necessity for defense methods for better adversarial robustness.
@@ 2.2 Adversarial Defense Methods
To mitigate the performance degradation caused by adversarial attacks, various adversarial defense methods have been developed. They can be grouped into three categories: training-based, ensembling-based, and denoising-based methods. Adversarial training improves the robustness of the model to adversarial examples through strategies like data augmentation (Si et al., 2021; Bao et al., 2021) and adversarial regularization (Madry et al., 2018; Zhu et al., 2020; Wang et al., 2021; Xi et al., 2022; Gao et al., 2023; Formento et al., 2024). However, adversarial training methods are limited as they assume similar train-test adversarial examples, and thus tend to overfit to specific types of adversarial attacks. Ensembling-based methods generate multiple variants of the input text at inference time and ensemble model predictions over all the variants (Ye et al., 2020; Zhou et al., 2021; Zeng et al., 2023; Li et al., 2023), but they can be inefficient given that model predictions are needed on every ensemble, increasing the inference time with the number of ensembles. More recently, denoising-based methods have been proposed to improve adversarial robustness by mapping the vector representation of the adversarial text to another point in the latent space that is close to the clean text (Nguyen Minh and Luu, 2022; Wang et al., 2023; Moon et al., 2023; Yuan et al., 2024; Ji et al., 2024). The denoised representation makes it more difficult to find vulnerable words to attack, thus improving adversarial robustness (Wang et al., 2023). Nevertheless, denoising might lead to very different representations of clean text and adversarial text, therefore changing the semantic meanings.
@@ 3.1 Training
The proposed diffusion defense model consists of a pretrained encoder for feature extraction, a transformer-based diffusion layer for noise prediction and reduction, and a classifier layer for output generation. The training process is split into two stages: adversarial training and diffusion training (Figure 1). The adversarial training stage employs a neural network-based adversarial training method like FreeLB++ (Li et al., 2021) or RSMI (Moon et al., 2023), which optimizes the encoder and classifier for robustness by perturbing the latent representation of the text input.
In the diffusion training stage, only the diffusion layer is trained to predict random noise added to the clean text hidden state at different timesteps, enabling it to denoise the adversarial hidden state at inference time. The pretrained encoder, however, is frozen during this stage. Since the pretrained encoder is only used for feature extraction, the diffusion training method is compatible with any neural network-based adversarial training method.
Given an input sequence of tokens , the pretrained encoder extracts the hidden state . A random Gaussian noise is sampled to perturb the hidden state . Sohl-Dickstein et al. (2015) define the forward diffusion process as a Markov Chain where at each timestep a Gaussian noise is sampled and added to the previous latent feature: , where , is the noisy hidden state at step and is a pre-calculated variance schedule changing with . As shown by Ho et al. (2020), this equation can be reformulated to calculate directly from by defining and , thus
At each training step, a random forward diffusion timestep is sampled from a uniform distribution. Therefore, the noisy hidden state is created from , , and . The diffusion layer consists of a time embedding and a transformer layer. The time embedding receives the diffusion timestep as input and produces an embedding , which is added to as input for the transformer layer. Finally, the transformer layer outputs the predicted noise , and mean square error (MSE) is used to compute the loss between the predicted noise and the actual sampled noise .
@@ 3.2 Inference
Leveraging the diffusion layer’s ability to predict noise at a given timestep , we utilize it as a denoiser during inference by iteratively performing the reverse diffusion steps, which sample from to produce the denoised hidden state
@@
Inference in DiffuseDef combines a one-step noising, a multi-step denoising, and an ensembling step. After the pretrained encoder extracts its hidden state , a set of Gaussian noise vectors are sampled to perform a single forward diffusion step. These noise vectors are then added to the hidden state following equation 1, resulting in a set of noisy hidden states , where denotes the number of denoising steps. The noisy hidden states are subsequently denoised through reverse diffusion steps, where noise is predicted by the diffusion layer and subtracted from the previous noisy hidden states. Unlike Ho et al. (2020), where the reverse diffusion step starts with pure noise sampled from a standard normal distribution, we assume the noisy hidden state is already an intermediate state in the reverse diffusion steps. This allows us to use a smaller number of than the training timestep to prevent the denoised hidden states from diverging substantially from the initial hidden state . This sequence of denoising steps creates the final denoised hidden states , which are averaged and used by the classifier to output the final predicted label. This process is summarized in Algorithm 1.
@@ 4 Experiments
Datasets.
We focus on two common NLP tasks in our experiments: topic classification and natural language inference (NLI). We conduct experiments on three common datasets: AG News (Zhang et al., 2015a), IMDB (Maas et al., 2011a), and QNLI datasets (Wang et al., 2018). We randomly split AGNews, IMDB, and QNLI datasets into train, validation, and test splits. Appendix A gives details on data preparation.
Evaluation.
Following previous work on adversarial defense, we use three benchmarking black-box attack methods to evaluate the robustness of DiffuseDef: TextFooler (TF) (Jin et al., 2020), TextBugger (TB) (Li et al., 2019), and Bert-Attack (BA) (Li et al., 2020). We also evaluated the robustness against two white-box attacks: T-PGD (Yuan et al., 2023) and SemAttack (Wang et al., 2022). Regarding evaluation metrics, we measure the clean accuracy (Clean%) on the test set, the accuracy under attack (AUA%), and the number of adversarial queries (#Query) needed for a successful attack. Higher scores on the three metrics denote a better robustness performance of a defense method. The accuracy on clean data is measured across the entire test set. The accuracy under attack and number of queries, due to the lengthy attacking process, is measured on a randomly sampled subset of 1000 examples from the test set. We use the TextAttack library as the adversarial evaluation framework. To ensure a fair comparison and high-quality adversarial examples, we follow the same evaluation constraints as in Li et al. (2021). The evaluation metrics are averaged based on experiments run with 5 random seeds.
@@ 4.1 Comparison to SOTA
We compare our proposed method with state-of-the-art adversarial defense approaches, trained using both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as backbones: Fine-tune: Fine-tuning pretrained models on downstream tasks with no defense method applied222“Fine-tuned” is a baseline approach used to illustrate the effect of adversarial attacks.. InfoBERT (Wang et al., 2021): Applying mutual-information-based regularizer during fine-tuning of pretrained models to improve robustness. FreeLB++ (Li et al., 2021): An adversarial training method improving on FreeLB(Zhu et al., 2020), which adds adversarial perturbations to word embedding during fine-tuning. EarlyRobust333We only run EarlyRobust with BERT as its implementation with RoBERTa has not been released. (Xi et al., 2022): Extracting early-bird subnetworks and pruning pretrained models for efficient adversarial training. RMLM (Wang et al., 2023): A denoising-based model combined with adversarial text detection. ATINTER (Gupta et al., 2023): A fine-tuned T5 model that rewrites the input adversarial texts, and the classifier predicts labels from the rewritten texts. RSMI (Moon et al., 2023): A two-stage training method that combines randomized smoothing and masked inference to improve adversarial robustness.